# From Open Vocabulary to Open World: Teaching Vision Language Models to Detect Novel Objects

### v3

@@ 4.3 Results
4.4 Qualitative Analysis
Fig. 5 shows a comparison between two SOTA methods and our method after completing the second task of M-OWODB and nu-OWODB. Our method provides more meaningful bounding boxes for unknown classes, without significantly losing performance on known classes, which makes it easier to make trade-off between precision and recall in various applications. We can see that although PROB [75] is able to detect a reasonable number of unknown objects, but there are also many high-confidence unknown bounding boxes associated with known objects. On the other hand, EO-OWODB [59] detects background as known objects while failing to detect several unknown objects. More discussion on the trade-off between precision and recall can be found in supplementary material.
@@ 4.5 Ablation Study
To understand the contribution of individual components, we disable some modules of our model to create a set of baseline models. Base Model is the vanilla Yolo World detector [5] prompted with known class names and a hand-crafted generic object name (“object”), in a zero shot manner. OWEL removes MSCAL modules and OOD scores, while the MSCAL replaces the pseudo unknown embedding $w_{\mathcal{U}}$ with the original generic prompt $w_{0}$.
From Tab. 5, we can see that OWEL significantly improves the U-Recall for unknown classes and mAP of known classes. On the other hand, OWEL increases absolute open set error. This indicates that optimizing known class embeddings with object detection loss not only learns embeddings most similar to image samples, but also learns some characteristics of this domain, which leads to more valid predictions and open set errors. MSCAL largely reduces the open set error, and achieves reasonable unknown recall without pseudo unknown embedding, because it detects unknown classes by correcting the open set error. OWEL and MSCAL complement each other to detect the largest proportion of unknown objects while reasonably maintaining the performance on known classes.
Additional ablation studies can be found in the supplemental material.
@@ 5 Conclusion
In this work, we propose a framework that enables open vocabulary models to operate in open world settings. We introduce Open World Embedding Learning (OWEL) and Multi-Scale Contrastive Anchor Learning (MSCAL) which enable the model to identify and incrementally learn unknown objects. We further propose a new benchmark to evaluate the performance of OWOD for autonomous driving. In the future, we will study open world object detection with various sensor modalities and data domains.
Supplementary Material
6 Additional Ablation Studies
6.1 Generic Objectness Prompt
Generic Prompt mAP $\uparrow$ U-Recall $\uparrow$ WI $\downarrow$ A-OSE $\downarrow$
object 71.9 16.6 0.0311 9070
entity 71.9 0.9 0.0353 10883
unknown 71.9 1.6 0.0361 10743
anything 71.9 5.9 0.0326 10154
everything 71.9 4.3 0.0334 10379
Table 6: Ablation study of different prompts on M-OWODB task 1, where 20 classes in PASCAL VOC are known.
We tried different prompts to estimate the embedding for generic objectness. Tab. 6 shows the ablation experiment on M-OWODB, which shows the word “object” is an appropriate. We define “object” as the generic prompt in one experiment, then reuse it in all OWOD benchmarks (M-OWODB, S-OWODB, and nu-OWODB) without hand-crafted selection of the prompt across domains. Though it is possible to estimate the generic prompt by designing a pretext task, but this will introduce bias towards known classes.
6.2 Fine-tuning YOLO-World
Method mAP $\uparrow$ U-Recall $\uparrow$ WI $\downarrow$ A-OSE $\downarrow$
YOLO-World (zero-shot) 21.8 2.1 0.0463 12316
YOLO-World (Fine-tuned) 30.0 4.9 0.0419 20039
Ours 28.1 45.5 0.0185 1724
Table 7: Open world performance on nu-OWODB task 1.
Although OVD models perform zero-shot detection by design, we can still fine-tune the model for better performance, especially on datasets from different domains. However, this introduces additional problems in the open-world scenario. As shown in Tab. 7, when we fine-tune the class embeddings of YOLO-World on nu-OWODB, the closed-set performance (mAP) improves, but the model’s resistance to OOD objects is significantly reduced, resulting in a higher A-OSE. In contrast, our method shows significant gains in both known and unknown performance, highlighting the importance of the proposed Pseudo Unknown Embedding and Multi-Scale Contrastive Anchor Learning.
6.3 Ablation on $\alpha$
As shown in Tab. 8, we try different $\alpha$ value varing from 0.2 to 0.8. The result shows that the choice of $\alpha$ does not make a significant impact on performance.
$\alpha$ mAP $\uparrow$ U-Recall $\uparrow$ WI $\downarrow$ A-OSE $\downarrow$
0.2 28.2 38.2 0.0183 1700
0.4 28.1 45.5 0.0185 1724
0.8 27.6 37.8 0.0166 1455
Table 8: Ablation study of $\alpha$ on nu-OWODB task 1.
7 Details of nu-OWODB
In this research, we present the nuScenes Open World Object Detection Benchmark (nu-OWODB), a novel benchmark designed to simulate the challenges of open-world object detection (OWOD) encountered in real world. Built on the nuImages subset of nuScenes [2], the benchmark encompasses 23 highly diverse and imbalanced object classes. The dataset is publicly available at www.nuscenes.org/nuimages.
We divide the classes into three major tasks: vehicles, pedestrians and obstacles. The task-category mapping relationship is shown in Tab. 9, and information about each task is provided in Tab. 10.
In addition, the naming convention in nuScenes is different from the natural language names of the classes. To fully utilize YOLO World’s zero shot capabilities, we define a natural language prompt for each class as shown in Tab. 11.
Table 9: Task-Category Mapping in nu-OWODB.
Task nuScenes Category
Task 1 - Vehicles vehicle.bicycle
vehicle.motorcycle
vehicle.car
vehicle.bus.bendy
vehicle.bus.rigid
vehicle.truck
vehicle.emergency.ambulance
vehicle.emergency.police
vehicle.construction
vehicle.trailer
Task 2 - Pedestrians human.pedestrian.adult
human.pedestrian.child
human.pedestrian.wheelchair
human.pedestrian.stroller
human.pedestrian.personal_mobility
human.pedestrian.police_officer
human.pedestrian.construction_worker
Task 3 - Obstacles movable_object.barrier
movable_object.trafficcone
movable_object.pushable_pullable
movable_object.debris
static_object.bicycle_rack
animal
Table 10: Task composition in nu-OWODB. The semantics of each task split and the number of associated training and test images and object instances are displayed.
Task IDs (→) Task 1 Task 2 Task 3
nu-OWODB Vehicles Pedestrians Obstacles
# classes 10 7 6
# training images 53850 34957 25682
# test images 13099 8473 6500
# training instances 274587 135870 147253
# test instances 64303 32710 39060
Table 11: Correspondence between class names and text prompts in nu-OWODB.
Class Name Text Prompt
vehicle.bicycle bicycle
vehicle.motorcycle motorcycle
vehicle.car car
vehicle.bus.bendy articulated bus
vehicle.bus.rigid rigid bus
vehicle.truck truck
vehicle.emergency.ambulance ambulance
vehicle.emergency.police police car
vehicle.construction construction vehicle
vehicle.trailer trailer
human.pedestrian.adult adult
human.pedestrian.child child
human.pedestrian.wheelchair wheelchair
human.pedestrian.stroller stroller
human.pedestrian.personal_mobility scooter
human.pedestrian.police_officer police officer
human.pedestrian.construction_worker construction worker
movable_object.barrier barrier
movable_object.trafficcone traffic cone
movable_object.pushable_pullable pushable and pullable object
movable_object.debris debris
static_object.bicycle_rack bicycle rack
animal animal
Figure 6: Visualization of indoor scenes on M-OWODB. We compare our method with PROB [75] and EO-OWOD [59] using their official checkpoints on M-OWODB task2. Face occlusions are added after model inference.
Figure 7: Reducing the threshold leads to a significant drop in precision. When EO-OWODB uses a threshold of 0.2, the number of bounding boxes of unknown classes increases slightly but also leads to many invalid bounding boxes.
8 Discussion on Evaluation Metrics
In this work, we use 4 commonly used evaluation metrics. For known classes, the commonly used metric is mean average precision (mAP). For unknown classes, the primary metric is unknown class recall (U-Recall), which assesses the ratio of detected unknown objects. In addition, we choose wilderness impact (WI) [6] and absolute open-set error (A-OSE) [41] as secondary evaluation metrics.
U-Recall and mAP are intuitive and widely adopted evaluation metrics in OWOD. A-OSE and WI are primarily designed for open set object detection, and evaluate the interference of unknown objects on the detection performance of known objects. A-OSE shows the absolute number of unknown objects detected as known classes at 0.5 IoU threshold. WI measures the impact of unknown objects on the model’s precision. The definition of WI is:
$WI=\frac{P_{\mathcal{K}}}{P_{\mathcal{K}\cup\mathcal{U}}}-1,$ (5)
where $P_{\mathcal{K}}$ is the precision in closed set and $P_{\mathcal{K}\cup\mathcal{U}}$ is the precision in open set. Following [19], we use the 0.8 recall threshold and 0.5 IoU threshold when calculating WI.
Although A-OSE and WI can somehow reflect the model’s confusion between known and unknown objects, some limitations remain. In Tab. 3, we can see that PROB [75] and EO-WOOD [59] experience catastrophic forgetting after learning the new class, but they achieve the best A-OSE and WI. As a more extreme example, if the model does not output any bounding box at all, A-OSE will be 0. As a result, A-OSE and WI are only meaningful when models have similar precision.
9 Additional Visualizations
In Fig. 5, we can see that our method provides high-quality bounding boxes for both known and unknown objects in outdoor scenes. Fig. 6 shows the detection results of indoor scenes on M-OWODB benchmark. Our method accurately detects objects in known classes and provides high-quality bounding boxes of unknown objects. PROB [75] provides a large amount of bounding boxes for unknown classes, but most of them do not capture objectness. On the other hand, many known objects are ignored. EO-OWOD [59], however, makes many classification errors within known classes, while the bounding boxes of unknown categories are few and closely associated with known categories (it actually correctly detects two persons, but the bounding boxes are overlapped with the unknown ones).
From Fig. 5 and Fig. 6 we can see that the detection confidence is not always meaningful in OWOD because bounding boxes with high confidence can also be invalid. Monotonically scaling up the confidence score does not change the detection performance. For all visualizations, we set the score threshold based on the principle of showing as many bounding boxes of unknown classes as possible without significantly reducing the precision of known classes. As shown in Fig. 7, a lower score threshold leads to higher recall for EO-OWODB, but it also leads to a collapse in precision. Although it is possible to set a threshold for each class separately, this will greatly reduce the ease of use of the model. The ease of use is ignored in evaluation metrics, but it is quite important for real-world applications.
10 Limitations and Future Work
Although our method achieves state-of-the-art performance in various benchmarks, some limitations still exist. For example, we assume object embeddings cluster around the word “object” in the semantic space, but there can also be outliers. Alternatively, we may construct more pseudo unknown embeddings, but this will increase computational overhead. For applications such as autonomous driving, we can further explore the use of LiDAR to detect class-agnostic objects and use the geometric information to perform 3D object detection in future work.

### v4

@@ 4.4 Ablation Study
To understand the contribution of individual components, we disable some modules of our model to create a set of baseline models. Base Model is the vanilla YOLO-World detector [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan] prompted with known class names and a hand-crafted generic object name (“object”), in a zero shot manner. OWEL removes MSCAL modules and OOD scores, while the MSCAL replaces the pseudo unknown embedding $w_{\mathcal{U}}$ with the original generic prompt $w_{0}$.
From Table˜5, we can see that OWEL significantly improves the U-Recall for unknown classes and mAP of known classes. On the other hand, OWEL increases absolute open set error. This indicates that optimizing known class embeddings with object detection loss not only learns embeddings most similar to image samples, but also learns some characteristics of this domain, which leads to more valid predictions and open set errors. MSCAL largely reduces the open set error, and achieves reasonable unknown recall without pseudo unknown embedding, because it detects unknown classes by correcting the open set error. OWEL and MSCAL complement each other to detect the largest proportion of unknown objects while reasonably maintaining the performance on known classes.
@@ 5 Conclusion
In this work, we propose a framework that enables open vocabulary object detectors to operate in open world settings without compromising their zero-shot capabilities. We introduce Open World Embedding Learning (OWEL) and Multi-Scale Contrastive Anchor Learning (MSCAL) which enable the model to identify and incrementally learn unknown objects. We further propose a new benchmark to evaluate the performance of OWOD for autonomous driving. In the future, we will study open world object detection with various sensor modalities and data domains.
Acknowledgments
The first two authors acknowledge the financial support from The University of Melbourne through the Melbourne Research Scholarship. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.